# Modeling Language Evolution Using a Spin Glass Approach

### v1

@@ Appendix A Dynamical roles of syntactic dependence classes
To clarify how implications constrain changes in the values of the syntactic parameters, it is useful to group them into classes according to their dependence depth — that is, the number of other parameters their relevance depends upon — labeling these groups $D_{0},D_{1},D_{2},\ldots,D_{6}$. For example, a parameter in $D_{3}$ is only relevant if 3 other parameters take specific values; otherwise it is undefined, which in our model is conflated with taking the default value -1.
Figure 10: Dynamical evolution of syntactic parameters grouped by dependence depth. (a) Temporal evolution of the normalized number of flipped parameters for each dependence class $D_{0},D_{1},\ldots,D_{6}$ in the deterministic limit ($\zeta\to\infty$). The 13 independent parameters in class $D_{0}$, which are unconstrained by implications, are the most active, flipping around half the times. (b) Same as (a), but with added weak symmetric random interactions ($\zeta=1000$, $\phi=0^{\circ}$). They suffice to freeze the system of parameters. (c) Same as (a), but with weak antisymmetric stochastic interactions ($\zeta=1000$, $\phi=90^{\circ}$), which keep the system fluid.
@@
First, we look at network dynamics in the deterministic limit ($\zeta\to\infty$), where dynamics are governed purely by implications, in the absence of the extra interactions, which in our model are randomly assigned. Figure 10a shows the temporal evolution of the normalized count of flipping parameters for each syntactic class. Independent parameters, the core set of 13 syntactic features in $D_{0}$ not dependent on implications, of course flip every second time step, on average. Changes in these free parameters propagate through the hierarchy, making parameters in $D_{1}$ those with the next highest probability to flip, then those in $D_{2}$, and so on. Ultimately, therefore, all syntactic changes are enabled, in this limit, by flipping the 13 independent parameters, out of the total 94, in the absence of random couplings. When the latter are introduced, even with minimal strength ($\zeta=1000$, see Fig. 10a and b for $\phi=0^{\circ}$ and $\phi=90^{\circ}$, respectively), the free parameters are not free anymore: for them, unconstrained by implications, the random couplings are the interactions there are, and they suffice to freeze the evolution of the free parameters in a few time steps, if symmetric or nearly symmetric. The other parameters then freeze as a result, see Fig. 10a for $\phi=0^{\circ}$. Beyond approximately $\phi=30^{\circ}$ the random interactions are asymmetric enough to keep the system fluid, and the fluidity is expressed in a similar hierarchy of flipping rates as the one seen in the absence of random couplings (as Fig. 10c shows for $\phi=90^{\circ}$).

### v2

@@ Appendix A Syntactic dependence
To clarify how implications constrain changes in the values of the syntactic parameters, we could proceed down the levels of Fig.A.1, but it is somewhat more perspicuous to group them into classes according to their dependence degree — that is, the number of other parameters their relevance depends upon — labeling these groups $D_{0},D_{1},D_{2},\ldots,D_{6}$. For example, a parameter in $D_{3}$ is only relevant if 3 other parameters take specific values; otherwise it is undefined, which in the first version of our model is conflated with taking the default value -1, in the second it is assigned value 0.
Figure A.1: Syntactic implications operate at different depth levels. The implications are shown as arrows pointing at the dependent parameters, numbered as in the PCM database (see [ceolin2021boundaries]; level $L_{k}$ with $k=0,\dots,10$ are arranged top to bottom, while the number $k^{\prime}$ of arrows impinging on each parameter correspond to its class $D_{k^{\prime}}$, with $k^{\prime}=0,\dots,6$. Boolean rules are not represented in this simple connectivity diagram, but are detailed in their Suppl. Table 2). Note that $L_{0}\equiv D_{0}$.
@@
First, we look at network dynamics in the deterministic limit ($\zeta\to\infty$), where dynamics are governed purely by implications, in the absence of the extra interactions, which in our model are randomly assigned. The model is the purely binary one. Figure A.2a shows the temporal evolution of the normalized count of flipping parameters for each syntactic class. Independent parameters, the core set of 13 syntactic features in $D_{0}$ not dependent on implications, of course flip every second time step, on average. Changes in these free parameters propagate through the hierarchy, making parameters in $D_{1}$ those with the next highest probability to flip, then those in $D_{2}$, and so on. Ultimately, therefore, all syntactic changes are enabled, in this limit, by flipping the 13 independent parameters, out of the total 94, in the absence of hidden couplings. When the latter are introduced, even with minimal strength ($\zeta=1000$, see Fig. A.2a and b for $\phi=0^{\circ}$ and $\phi=90^{\circ}$, respectively), the free parameters are not free anymore: for them, unconstrained by implications, the hidden couplings are the interactions there are, and they suffice to freeze the evolution of the free parameters in a few time steps, if symmetric or nearly symmetric. The other parameters then freeze as a result, see Fig. A.2a for $\phi=0^{\circ}$. Beyond approximately $\phi=30^{\circ}$ the hidden interactions are asymmetric enough to keep the system fluid, and the fluidity is expressed in a similar hierarchy of flipping rates as the one seen in the absence of hidden couplings (as Fig. A.2c shows for $\phi=90^{\circ}$).
Figure A.2: Dynamical evolution of syntactic parameters grouped by dependence number. (a) Temporal evolution of the normalized number of flipped parameters for each class $D_{0},D_{1},\ldots,D_{6}$ in the deterministic limit ($\zeta\to\infty$), in the first version of the model. The 13 independent parameters in class $D_{0}$, which are unconstrained by implications, are the most active, flipping around half the times. (b) Same as (a), but with added weak symmetric hidden interactions ($\zeta=1000$, $\phi=0^{\circ}$). They suffice to freeze the system of parameters. (c) Same as (a), but with weak antisymmetric stochastic interactions ($\zeta=1000$, $\phi=90^{\circ}$), which keep the system fluid.